# The Picard group of the Baily-Borel compactification of the moduli space of polarized K3 surfaces

### v1

@@ 3.1 Bruinier-Stein’s vector-valued Hecke operator
$\widetilde{\mathbf{Y}}_{\alpha^{2}}:={\rm Mp}_{2}(\mathbb{Z})\cdot\widetilde{\mathbf{g}}_{\alpha}\cdot{\rm Mp}_{2}(\mathbb{Z}).$
For any element $\tilde{g}\cdot\widetilde{\mathbf{g}}_{\alpha}\cdot\tilde{g}^{\prime}\in\widetilde{\mathbf{Y}}_{\alpha^{2}}$, define the action
$\mathfrak{e}_{\lambda}\mid[\tilde{g}\cdot\widetilde{\mathbf{g}}_{\alpha}\cdot\tilde{g}^{\prime}]:=\mathfrak{e}_{\lambda}\mid_{G}[\tilde{g}]\cdot\mid_{G}[\widetilde{\mathbf{g}}_{\alpha}]\cdot\mid_{G}[\tilde{g}^{\prime}],$
where $\mathfrak{e}_{\gamma}\mid_{G}[\tilde{g}]=\rho_{G}(\tilde{g})^{-1}(\mathfrak{e}_{\gamma})$ is the Weil representation and $\mathfrak{e}_{\gamma}\mid_{G}[\widetilde{\mathbf{g}}_{\alpha}]=\mathfrak{e}_{\alpha\gamma}$. This action is proved to be well-defined in [11, Proposition 5.1].
@@
It is easy to see that this extension can be made also in the case when $\alpha=0$. This further extension will be used in Section 4.
@@
Definition 3.2.
[11, Definition 5.5] Let $\alpha>0$ be a positive integer. Express $\widetilde{\mathbf{Y}}_{\alpha^{2}}$ as a disjoint union of left cosets
@@
$\mathbf{T}_{\alpha^{2}}(f):=\alpha^{k-2}\sum_{i}\sum_{\gamma\in G}(f_{\gamma}\mid_{k}[\tilde{\delta}_{i}])\otimes(\mathfrak{e}_{\gamma}\mid[\tilde{\delta}_{i}]),$
@@
This definition is proved to be well-defined in [11, Section 5]. The main properties of these Hecke operators are summarized in the following proposition.
Proposition 3.3.
[11, Theorem 4.12, Theorem 5.6] The Hecke operators defined in Definition 3.2 satisfy the following properties.
  • These Hecke operators take cusp forms to cusp forms. They are self-adjoint with respect to the Petersson inner product, and the Hecke operators $\{\mathbf{T}_{\alpha^{2}}:\gcd(\alpha,N)=1\}$ generate a commutative subalgebra of $\operatorname{End}_{\mathbb{C}}({\mathrm{Cusp}}_{k}(\rho_{G}))$.
@@ 3.2 Non-vanishing of $L$-values
$L(f,s):=\sum_{\begin{subarray}{c}\alpha\geq 1\\ \gcd(\alpha,N)=1\end{subarray}}\frac{\lambda(\alpha^{2})}{\alpha^{s}}.$ (3.1)
@@
The case when $k\in\mathbb{Z}$ is [25, Proposition 3.3] and the result is slightly stronger. Now we consider the case when $k\in\mathbb{Z}+\frac{1}{2}$, then $\mathrm{sign}(G)$ is odd. We will establish the theorem by comparing Bruinier-Stein’s Hecke operators with the scalar-valued Hecke operators defined by Shimura [31].
  1. $(i)$
    Since $\mathrm{sign}(G)$ is odd, the oddity formula [13, p. 383 (30)] implies that $4\mid N$. There is a map
@@ 1. $(i)$
    where $\epsilon_{d}=1$ if $d\equiv 1\mod 4$, $\epsilon_{d}=i$ if $d\equiv-1\mod 4$, and $\left(\frac{\cdot}{\cdot}\right)$ is the Jacobi symbol. This map lifts to ${\rm Mp}_{2}(\mathbb{Z})$ when restricted to $\Gamma_{1}(4)$. We denote by $\Delta(N)$ the image of any principal congruence subgroup $\Gamma(N)\subset\Gamma_{1}(4)$ under this lift. For a character $\chi$ of $(\mathbb{Z}/N\mathbb{Z})^{\times}$ and $\gamma\in G$, define the element
    $v_{\gamma,\chi}:=\sum\limits_{d\in(\mathbb{Z}/N\mathbb{Z})^{\times}}\chi(d)^{-1}\mathfrak{e}_{d\gamma}.$
    Since the Weil representation is trivial on $\Delta(N)$, the function $F_{\gamma,\chi}:=\langle f,v_{\gamma,\chi}\rangle$ is then an element in ${\mathrm{Cusp}}_{k}(\Delta(N))$, where ${\mathrm{Cusp}}_{k}(\Delta(N))$ is the space of scalar-valued cusp forms of weight $k$ and level $\Delta(N)$ (cf. [31]).
  2. $(ii)$
@@
    by $\begin{pmatrix}\alpha^{2}&0\\ 0&1\end{pmatrix}\mapsto\widetilde{\mathbf{g}}_{\alpha}$. For every prime $p$ satisfying $\gcd(p,N)=1$, choose an element $R_{p^{2n-a}}\in{\rm SL}_{2}(\mathbb{Z})$ as a lift of the element $\begin{pmatrix}p^{-2n+a}&0\\ 0&p^{2n-a}\end{pmatrix}\in{\rm SL}_{2}(\mathbb{Z}/N\mathbb{Z})$ and consider the set
    $\Sigma_{p^{2n}}:=\left\{\kappa\left(R_{p^{2n-a}}\begin{pmatrix}p^{2n-a}&bN\\ 0&p^{a}\end{pmatrix}\right)\mid 0\leq a\leq 2n,~{}0\leq b<p^{a},~{}\gcd(b,p^{\min\{a,2n-a\}})=1\right\}.$
    Since $R_{p^{2n-a}}\begin{pmatrix}p^{2n-a}&bN\\ 0&p^{a}\end{pmatrix}\equiv\begin{pmatrix}1&0\\ 0&p^{2n}\end{pmatrix}\mod N$, the element $\kappa\left(R_{p^{2n-a}}\begin{pmatrix}p^{2n-a}&bN\\ 0&p^{a}\end{pmatrix}\right)$ lies in the metaplectic double cover of ${\rm GL}_{2}(\mathbb{R})$. As shown in [35, §4], this set can be served as a system of representatives for
    ${\rm Mp}_{2}(\mathbb{Z})\backslash{\rm Mp}_{2}(\mathbb{Z})\Big{(}\left(\begin{array}[]{cc}p^{2n}&0\\ 0&1\end{array}\right),1\Big{)}{\rm Mp}_{2}(\mathbb{Z})\cong\Delta(N)\backslash\Delta(N)\Big{(}\left(\begin{array}[]{cc}1&0\\ 0&p^{2n}\end{array}\right),p^{n}\Big{)}\Delta(N).$
  3. $(iii)$
    The advantage of choosing $\Sigma_{p^{2n}}$ as coset representatives is that the actions of these representatives via the Weil representation are particularly simple. Let $\tilde{g}=\Big{(}g=\begin{pmatrix}a&b\\ c&d\end{pmatrix},\phi_{g}(\tau)\Big{)}\in{\rm Mp}_{2}(\mathbb{Z})$ with $N\mid b,~{}N\mid c$ For $\gamma\in G$, we have $\rho_{G}(\tilde{g})\mathfrak{e}_{\gamma}=\chi_{G}(\tilde{g})\mathfrak{e}_{d\gamma}$ for an explicit character $\chi_{G}$ (cf. [7, Theorem 5.4]). It follows that for every $\tilde{g}\in\Sigma_{p^{2n}}$, we have
@@ 3. $(iii)$
  4. $(iv)$
    In [31], Shimura introduced a Hecke operator
@@ 4. $(iv)$
    By the discussions in $(ii)$, it is easy to see that
@@
    Now we prove the claim. First observe that $F_{\gamma,\chi}(N\cdot)\in{\mathrm{Cusp}}_{k}(\Delta_{0}(N^{2}),\psi)$ for some character $\psi$ of $(\mathbb{Z}/N^{2}\mathbb{Z})^{\times}$. Furthermore, we have $(\mathcal{T}_{\alpha^{2}}^{\Delta(N)}F_{\gamma,\chi})(N\cdot)=\mathcal{T}_{\alpha^{2}}^{\Delta_{0}(N^{2}),\psi}(F_{\gamma,\chi}(N\cdot))$. This is an analog of the classical way of treating a modular form of level $\Gamma(N)$ as a modular form of level $\Gamma_{0}(N^{2})$ with a Nebentypus. Now let $p$ be a prime such that $\gcd(p,N)=1$. By the theory of Shimura lifts, $\chi(p)\lambda(p^{2})$ is the $p$-th Fourier coefficient of a normalized cuspidal Hecke eigenform of weight $12k-$ 1. Hence the Ramanujan conjecture proved by Deligne implies that
@@
    Since $p\geq 3$, by the explicit algebraic relations between the Hecke operators $\{\mathcal{T}_{p^{2n}}^{\Delta_{0}(N^{2}),\psi}\mid n\geq 0\}$ in [29, Theorem 1.1], one can inductively prove that
@@
3.3 The case when $p\mid N$
The Hecke operators $\{\mathbf{T}_{\alpha^{2}}:\gcd(\alpha,N)\neq 1\}$ are trickier since the Hecke algebra generated by $\{\mathbf{T}_{p^{2n}}:n\geq 1\}$ is in general non-commutative if $p\mid N$ (cf. [11, Remark 5.7]). However, regarding our goals, a comparison result of these Hecke operators with certain scalar-valued Hecke operators is sufficient.
@@ 3.3 The case when $p\mid N$
  • when $p=2$, there exists $\mu\in G$, $02\mu=$ 0, such that $2\mathfrak{q}(\mu)+(\mu,\gamma)\neq 0\mod\mathbb{Z}$.
@@
$\mathrm{T}^{\mathfrak{q}(\gamma)}_{p^{2n}}(F):=p^{(k-2)n}\sum_{b=0}^{p^{2n}-1}\mathbf{e}(-b\mathfrak{q}(\gamma))F\mid_{k}\left[\Big{(}\left(\begin{array}[]{cc}1&b\\ 0&p^{2n}\end{array}\right),p^{n}\Big{)}\right]$
@@
$\widetilde{\mathbf{Y}}_{p^{2n}}={\rm Mp}_{2}(\mathbb{Z})\cdot\Big{(}\begin{pmatrix}p^{2n-a}&b\\ 0&p^{a}\end{pmatrix},\sqrt{p^{a}}\Big{)},$
where $0\leq a\leq 2n,~{}0\leq b<p^{a}$ and ${}\gcd(b,p^{\min\{a,2n-a\}})=1$ (cf. [32, §5.1]). Set $\tilde{\delta}_{a,b}=\Big{(}\begin{pmatrix}p^{2n-a}&b\\ 0&p^{a}\end{pmatrix},\sqrt{p^{a}}\Big{)}$. Then we have
@@
for every $a<2n$. The action of $\tilde{\delta}_{a,b}$ is well-understood. For instance, if $p$ is odd, there is an explicit formula of $\mathfrak{e}_{\mu}\mid[\tilde{\delta}_{a,b}]$ proved in [32, Theorem 5.2] and [8, Proposition 5.3]. In general, it is of the form
@@
with $\lambda\in G^{p}$ if $p$ is odd, and $\lambda=2m\mu+\beta$ for some $\beta\in G^{2\ast}$ and $m\in\mathbb{Z}$ if $p=2$ (cf. [30, Theorem 4.7] and [33, Theorem 1]). As $\gamma$ satisfies the condition $(\ast_{p})$, then $\langle\mathfrak{e}_{\mu}\mid[\widetilde{\delta}_{a,b}],\mathfrak{e}_{\gamma}\rangle=0$ for every $a<2n$. ∎
@@
Proposition 3.7.
[25, Corollary 3.5] Fix a finite set $S$ of primes. Let $f\in{\mathrm{Mod}}_{k}(\rho_{G})$ and let $\gamma,\mu\in G$ satisfy the following conditions:
@@
$v_{\gamma,S}^{\mu}:=\sum_{I\subset S}(-1)^{|I|}\mathfrak{e}_{\gamma_{I}^{\mu}},$
where $\gamma_{I}^{\mu}\in G$ is the element whose $p$-adic component is equal to that of $\mu$ when $p\in I$ and other $p$-adic components remain the same as those of $\gamma$. Then if $\prod\limits_{p\in S}\sum\limits_{n=1}^{\infty}\frac{\mathbf{T}_{p^{2n}}}{p^{ns}}(f)=0$, $\langle f,v^{\mu}_{\gamma,S}\rangle=0$.
@@
4 Modular forms as Theta series
This section studies the space of cusp forms ${\mathrm{Cusp}}_{k}(\rho_{M})$ with respect to an even negative definite lattice $M$. The goal is to show that under certain conditions of $\rho_{M}$, the space ${\mathrm{Cusp}}_{k}(\rho_{M})$ is spanned by the Theta series. Throughout this section, $M$ is an even negative definite lattice of rank $r$.
4.1 Projection to ${\mathrm{Cusp}}^{\theta}_{k}(\rho_{M})$
In [16], Eichler and Zagier introduced a differential operator transforming a $C^{\infty}$ function on $\mathbb{H}_{2}$ into a $C^{\infty}$ function on $\mathbb{H}\times\mathbb{H}$. To define this, we need Gegenbauer polynomials $G^{h}_{r}(x,y)$, which are defined by

### v2

@@ 3.1. Bruinier–Stein’s vector-valued Hecke operator
$\widetilde{\mathbf{Y}}_{\alpha^{2}}\coloneqq{\rm Mp}_{2}(\mathbb{Z})\cdot\widetilde{\mathbf{g}}_{\alpha}\cdot{\rm Mp}_{2}(\mathbb{Z}).$
For any element $\tilde{g}\cdot\widetilde{\mathbf{g}}_{\alpha}\cdot\tilde{g}^{\prime}\in\widetilde{\mathbf{Y}}_{\alpha^{2}}$, the action is defined by:
$\mathfrak{e}_{\lambda}\mid[\tilde{g}\cdot\widetilde{\mathbf{g}}_{\alpha}\cdot\tilde{g}^{\prime}]\coloneqq\mathfrak{e}_{\lambda}\mid_{G}[\tilde{g}]\cdot\mid_{G}[\widetilde{\mathbf{g}}_{\alpha}]\cdot\mid_{G}[\tilde{g}^{\prime}],$
where $\mathfrak{e}_{\gamma}\mid_{G}[\tilde{g}]=\rho_{G}(\tilde{g})^{-1}(\mathfrak{e}_{\gamma})$ is the Weil representation and $\mathfrak{e}_{\gamma}\mid_{G}[\widetilde{\mathbf{g}}_{\alpha}]=\mathfrak{e}_{\alpha\gamma}$. This action is well-defined by [BS, Proposition 5.1].
@@
It is easy to see that this extension can also be made in the case when $\alpha=0$. This further extension will be used in Section 4.
@@
Definition 3.2 ([BS, Def. 5.5]).
Let $\alpha>0$ be a positive integer. Express $\widetilde{\mathbf{Y}}_{\alpha^{2}}$ as a disjoint union of left cosets
@@
$\mathbf{T}_{\alpha^{2}}(f)\coloneqq\alpha^{k-2}\sum_{i}\sum_{\gamma\in G}(f_{\gamma}\mid_{k}[\tilde{\delta}_{i}])\otimes(\mathfrak{e}_{\gamma}\mid[\tilde{\delta}_{i}]),$
@@
This definition is proved to be well-defined in [BS, Section 5]. The main properties of these Hecke operators are summarized in the following proposition.
Proposition 3.3 ([BS, Thm. 4.12, Thm 5.6]).
The Hecke operators defined in Definition 3.2 satisfy the following properties:
  • These Hecke operators take cusp forms to cusp forms.
  • They are self-adjoint with respect to the Petersson inner product, and the Hecke operators
    $\{\mathbf{T}_{\alpha^{2}}:\gcd(\alpha,N)=1\}$
    generate a commutative subalgebra of $\operatorname{End}_{\mathbb{C}}({\mathrm{Cusp}}_{k}(\rho_{G}))$.
@@ 3.2. Non-vanishing of $L$-values
$L(f,s)\coloneqq\sum_{\begin{subarray}{c}\alpha\geq 1\\ \gcd(\alpha,N)=1\end{subarray}}\frac{\lambda(\alpha^{2})}{\alpha^{s}}.$ (3.1)
@@
The case when $k\in\mathbb{Z}$ is [Ma24, Proposition 3.3] and the result is slightly stronger. Now we consider the case when $k\in\mathbb{Z}+\frac{1}{2}$, then $\mathrm{sign}(G)$ is odd. We will establish the theorem by comparing Bruinier–Stein’s Hecke operators with the scalar-valued Hecke operators defined by Shimura [Shimura73].
Since $\mathrm{sign}(G)$ is odd, the oddity formula [CS87, p. 383 (30)] implies that $4\mid N$. There is a map
@@
where $\epsilon_{d}=1$ if $d\equiv 1\bmod 4$, $\epsilon_{d}=i$ if $d\equiv-1\bmod 4$, and $\left(\frac{\cdot}{\cdot}\right)$ is the Jacobi symbol. This map lifts to ${\rm Mp}_{2}(\mathbb{Z})$ when restricted to $\Gamma_{1}(4)$. We denote by $\Delta(N)$ the image of any principal congruence subgroup $\Gamma(N)\subset\Gamma_{1}(4)$ under this lift. For a character $\chi$ of $(\mathbb{Z}/N\mathbb{Z})^{\times}$ and $\gamma\in G$, define the element
$v_{\gamma,\chi}\coloneqq\sum\limits_{d\in(\mathbb{Z}/N\mathbb{Z})^{\times}}\chi(d)^{-1}\mathfrak{e}_{d\gamma}.$
Since the Weil representation is trivial on $\Delta(N)$, the function $F_{\gamma,\chi}\coloneqq\langle f,v_{\gamma,\chi}\rangle$ is then an element in ${\mathrm{Cusp}}_{k}(\Delta(N))$, where ${\mathrm{Cusp}}_{k}(\Delta(N))$ is the space of scalar-valued cusp forms of weight $k$ and level $\Delta(N)$ (cf. [Shimura73]).
@@
by $\left(\begin{smallmatrix}\alpha^{2}&0\\ 0&1\end{smallmatrix}\right)\mapsto\widetilde{\mathbf{g}}_{\alpha}$. For every prime $p$ satisfying $\gcd(p,N)=1$, choose an element $R_{p^{2n-a}}\in{\rm SL}_{2}(\mathbb{Z})$ as a lift of the element $\left(\begin{smallmatrix}p^{-2n+a}&0\\ 0&p^{2n-a}\end{smallmatrix}\right)\in{\rm SL}_{2}(\mathbb{Z}/N\mathbb{Z})$ and consider the set
$\Sigma_{p^{2n}}\coloneqq\left\{\kappa\left(R_{p^{2n-a}}\begin{pmatrix}p^{2n-a}&bN\\ 0&p^{a}\end{pmatrix}\right)\mid 0\leq a\leq 2n\right.,\left.0\leq b<p^{a},~{}\gcd(b,p^{\min\{a,2n-a\}})=1\right\}.$
Since the matrix $R_{p^{2n-a}}\left(\begin{smallmatrix}p^{2n-a}&bN\\ 0&p^{a}\end{smallmatrix}\right)$ is congruent to $\left(\begin{smallmatrix}1&0\\ 0&p^{2n}\end{smallmatrix}\right)$ modulo $N$, its image under $\kappa$ lies in the metaplectic double cover of ${\rm GL}_{2}(\mathbb{R})$. As shown in [Wr14, §4], this set can serve as a system of representatives for
${\rm Mp}_{2}(\mathbb{Z})\backslash{\rm Mp}_{2}(\mathbb{Z})\left(\begin{pmatrix}p^{2n}&0\\ 0&1\end{pmatrix},1\right){\rm Mp}_{2}(\mathbb{Z})\cong\Delta(N)\backslash\Delta(N)\left(\begin{pmatrix}1&0\\ 0&p^{2n}\end{pmatrix},p^{n}\right)\Delta(N).$
The advantage of choosing $\Sigma_{p^{2n}}$ as coset representatives is that the actions of these representatives via the Weil representation are particularly simple. Let $\tilde{g}=\left(g=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right),\phi_{g}(\tau)\right)\in{\rm Mp}_{2}(\mathbb{Z})$ with $N\mid b,~{}N\mid c$ For $\gamma\in G$, we have $\rho_{G}(\tilde{g})\mathfrak{e}_{\gamma}=\chi_{G}(\tilde{g})\mathfrak{e}_{d\gamma}$ for an explicit character $\chi_{G}$ (cf. [Bo00, Thm. 5.4]). It follows that for every $\tilde{g}\in\Sigma_{p^{2n}}$, we have
@@
In [Shimura73], Shimura introduced a Hecke operator
@@
By the discussions above, it is easy to see that
@@
Now we prove the claim. First observe that $F_{\gamma,\chi}(N\cdot)\in{\mathrm{Cusp}}_{k}(\Delta_{0}(N^{2}),\psi)$ for some character $\psi$ of $(\mathbb{Z}/N^{2}\mathbb{Z})^{\times}$. Furthermore, we have
$(\mathcal{T}_{\alpha^{2}}^{\Delta(N)}F_{\gamma,\chi})(N\cdot)=\mathcal{T}_{\alpha^{2}}^{\Delta_{0}(N^{2}),\psi}(F_{\gamma,\chi}(N\cdot)).$
This is an analog of the classical way of treating a modular form of level $\Gamma(N)$ as a modular form of level $\Gamma_{0}(N^{2})$ with a Nebentypus. Now let $p$ be a prime such that $\gcd(p,N)=1$. By the theory of Shimura lifts, $\chi(p)\lambda(p^{2})$ is the $p$-th Fourier coefficient of a normalized cuspidal Hecke eigenform of weight $12k-$ 1. Hence the Ramanujan conjecture proved by Deligne implies that
@@
Since $p\geq 3$, by the explicit algebraic relations between the Hecke operators $\{\mathcal{T}_{p^{2n}}^{\Delta_{0}(N^{2}),\psi}\mid n\geq 0\}$ in [Pu14, Thm. 1.1], one can inductively prove that
@@
3.3. The case when $p\mid N$
The Hecke operators $\{\mathbf{T}_{\alpha^{2}}:\gcd(\alpha,N)\neq 1\}$ are trickier since the Hecke algebra generated by $\{\mathbf{T}_{p^{2n}}:n\geq 1\}$ is in general non-commutative if $p\mid N$ (cf. [BS, Remark 5.7]). However, regarding our goals, a comparison result of these Hecke operators with certain scalar-valued Hecke operators is sufficient.
@@ 3.3. The case when $p\mid N$
  • when $p=2$, there exists $\mu\in G$, $02\mu=$ 0, such that $2\mathfrak{q}(\mu)+(\mu,\gamma)\neq 0\bmod\mathbb{Z}$.
@@
$\mathrm{T}^{\mathfrak{q}(\gamma)}_{p^{2n}}(F)\coloneqq p^{(k-2)n}\sum_{b=0}^{p^{2n}-1}\mathbf{e}(-b\mathfrak{q}(\gamma))F\mid_{k}\left[\Big{(}\left(\begin{array}[]{cc}1&b\\ 0&p^{2n}\end{array}\right),p^{n}\Big{)}\right]$
@@
$\widetilde{\mathbf{Y}}_{p^{2n}}=\bigcup_{a,b}{\rm Mp}_{2}(\mathbb{Z})\cdot\Big{(}\begin{pmatrix}p^{2n-a}&b\\ 0&p^{a}\end{pmatrix},\sqrt{p^{a}}\Big{)},$
where $0\leq a\leq 2n,~{}0\leq b<p^{a}$ and ${}\gcd(b,p^{\min\{a,2n-a\}})=1$ (cf. [St15, §5.1]). Set $\tilde{\delta}_{a,b}=\left(\left(\begin{smallmatrix}p^{2n-a}&b\\ 0&p^{a}\end{smallmatrix}\right),\sqrt{p^{a}}\right)$. Then we have
@@
for every $a<2n$. The action of $\tilde{\delta}_{a,b}$ is well-understood. For instance, if $p$ is odd, there is an explicit formula of $\mathfrak{e}_{\mu}\mid[\tilde{\delta}_{a,b}]$ proved in [St15, Thm. 5.2] and [BCJ19, Proposition 5.3]. In general, it is of the form
@@
with $\lambda\in G^{p}$ if $p$ is odd, and $\lambda\in G^{2}\cup G^{2\ast}$ if $p=2$ (cf. [Sc09, Thm. 4.7] and [Str13, Thm. 1]). As $\gamma$ satisfies the condition $(\ast_{p})$, it follows that $\langle\mathfrak{e}_{\mu}\mid[\widetilde{\delta}_{a,b}],\mathfrak{e}_{\gamma}\rangle=0$ for every $a<2n$. ∎
@@
Proposition 3.7 ([Ma24, Cor. 3.5]).
Fix a finite set $S$ of primes. Let $f\in{\mathrm{Mod}}_{k}(\rho_{G})$ and let $\gamma,\mu\in G$ satisfy the following conditions:
@@
$v_{\gamma,S}^{\mu}\coloneqq\sum_{I\subset S}(-1)^{|I|}\mathfrak{e}_{\gamma_{I}^{\mu}},$
where $\gamma_{I}^{\mu}\in G$ is the element whose $p$-adic component is equal to that of $\mu$ when $p\in I$ and other $p$-adic components remain the same as those of $\gamma$. Then $\prod\limits_{p\in S}\sum\limits_{n=1}^{\infty}\frac{\mathbf{T}_{p^{2n}}}{p^{ns}}(f)=0$ implies that $\langle f,v^{\mu}_{\gamma,S}\rangle=0$.
@@
4. Modular forms as theta series
This section studies the space of cusp forms ${\mathrm{Cusp}}_{k}(\rho_{M})$ with respect to a positive definite even lattice $M$. Our goal is to show that under certain specific conditions on $\rho_{M}$, the space ${\mathrm{Cusp}}_{k}(\rho_{M})$ is spanned by the theta series. Throughout this section, $M$ denotes a positive definite even lattice of rank $r$.
4.1. Projection to ${\mathrm{Cusp}}^{\theta}_{k}(\rho_{M})$
In [EZ], Eichler and Zagier introduced a differential operator transforming a $C^{\infty}$ function on $\mathbb{H}_{2}$ into a $C^{\infty}$ function on $\mathbb{H}\times\mathbb{H}$. To define this operator, we need Gegenbauer polynomials $G^{h}_{r}(x,y)$, which are defined by the generating function: